OpenStudy (anonymous):

Simplify: (3x − 5) + (3x + 6)

OpenStudy (anonymous):

6x + 1

OpenStudy (anonymous):

you add 3x and 3x together then -5 and 6 together .. 6x + 1 understand?

OpenStudy (anonymous):

Thanks This one was tuff for me

OpenStudy (anonymous):

Add like terms and simplify. 3x + 3x = 6x -5 + 6 = 1 Final answer: 6x + 1

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Hunter: Try this one for more practice. (4x-7) + (12x+5) =